OpenStudy (anonymous):

X+2=3x-6-6x How to slope this??

OpenStudy (anonymous):

I mean solve

OpenStudy (anonymous):

move the x+2 over to the other side so that you have 0=3x-6-6x-x-2

OpenStudy (anonymous):

x+2=-3x-6 4x=-8 x=-2

OpenStudy (anonymous):

collect like terms

OpenStudy (anonymous):

ok, you could have left the x

OpenStudy (anonymous):

anyway, then bring the x with the coefficient, over to the other side and solve

OpenStudy (anonymous):

damn this is a crappy explanation

OpenStudy (anonymous):

I got 0=3x-4-5x

OpenStudy (anonymous):

let me do this properly for you

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[x+2= 3x-6-6x\] \[x=3x-6-6x-2\] \[x-3x+6x=-6-2\] \[4x=-8\] \[\left(\begin{matrix}4x \\ 4\end{matrix}\right)=\left(\begin{matrix}-8 \\ 4\end{matrix}\right)\] \[x=-2\]

OpenStudy (anonymous):

so here is what I got X+2=3x-6-6x x+2=-3x-6 X=-3x-4 -4x=-4/-4 x=1

OpenStudy (anonymous):

so i moved all the x's to one side and tall the other numbers to the other then simplified both sides and divided the 4 out of the -8 to be left with only 1 x

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

when you move something from one side to the other it changes signs so 2x= 3 move the 2x over and it is 0=-2x+3

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

I got it thanks, My internet shutted down, thankyou again